OpenStudy (anonymous):

State whether the ratios form a proportion 25/35 and 35/45 A. yes B. no

OpenStudy (anonymous):

@cheyennerae1605 and @Vocaloid and

Vocaloid (vocaloid):

basically, you want to find out if 25/35 = 35/45 if they're equal, then it's a proportion if they're not equal, then it's not a proportion

OpenStudy (anonymous):

I know that but how would u find out that's its equal or not

Vocaloid (vocaloid):

you can use a calculator or simplify them by hand

Vocaloid (vocaloid):

25/35 = (5*5)/(5*7) = 5/7 35/45 = (5*7)/(5*9) = 7/9 so they're not equal

OpenStudy (anonymous):

ok can u help me with a few more Q?

Vocaloid (vocaloid):

actually, I need to get going, sorry about that D:

OpenStudy (anonymous):

make a new post ddaddy and i will try to help you

OpenStudy (anonymous):

Solve for the variable 3/9=p/24 A.1/8 B.1/3 C. 3 D.8 thx @nignog9000

OpenStudy (anonymous):

D

OpenStudy (anonymous):

ok can u help with more?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Solve for the variable k/3=16/24 A.1/2 B.2/3 C.2 D.6

OpenStudy (anonymous):

C

OpenStudy (anonymous):

The ratio of boys to girls in a youth class is 3 to 4.There are 16 girls in the chorus. how many boys are in the chorus? A.6 B.9 C.12 D.15

OpenStudy (anonymous):

C

OpenStudy (anonymous):

If i have help it would be nice for you to give me a medal xD

OpenStudy (anonymous):

What is the value of x? x/6=15/18 A.3 B.5 C.45 D.90 -sorry bout that hold on

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

ok thx so much if I need more help ill tag u if that's fine?

OpenStudy (anonymous):

yea any time

OpenStudy (anonymous):

Find the scale factor A.3 B.1/3 C.9/6 D.3/2 @nignog9000

OpenStudy (anonymous):

A

OpenStudy (anonymous):

It might be B I am not positive on this one

OpenStudy (anonymous):

Rectangles PQRS and TUVW are similar. Find the length of UV. A. 14mm B. 15mm C. 21mm D.31mm

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Triangles GHI and JKL are similar. What is the length of JL? A.12ft B.18ft C.21ft D.27ft

OpenStudy (anonymous):

B

OpenStudy (anonymous):

On a map represents 50 km. Two cities are 3.5 cm apart on the map. How far apart are the actual cities? A.53.5km B.150.5km C.155km D.175km

OpenStudy (anonymous):

on a map , 1cm represents 50km sorry forgot that part

OpenStudy (anonymous):

D

OpenStudy (anonymous):

what is the scale factor for the similar figures? A.1/6 B.1/3 C.2/3 D.3

OpenStudy (anonymous):

I am not good at scale factor it makes my brain hurt xD if you have a specific question about this problem I might be able to help

OpenStudy (anonymous):

what answer could be equal to the figures? @vbroberson

OpenStudy (anonymous):

@Ddaddysgirl13 I'm not sure...